# Sp2360: Sparse-view 360 Scene Reconstruction using Cascaded 2D Diffusion Priors

### v1

@@ 3.1 Sp2360 for Sparse-View 3D Reconstruction
For training 3DGS in an iterative fashion, as outlined above, special precaution has to be taken to prevent overfitting on the initial views. An optimal schedule involves finding the number of iterations per cycle $k$ and the hyperparameters for the original 3DGS [30]. We provide a detailed evaluation in App. A.3. In the following, we will detail the in-painting and artifact removal stages of our pipeline.
@@ 3.3 Removing Sparse-View Artifacts
We fine-tune an image-conditioned diffusion model [4] to edit images based on short, user-friendly edit instructions. Here, the UNet $\boldsymbol{\epsilon_{\theta}}$ is trained to predict noise in $z_{t}$ conditioned on a given image $c_{I}$, in addition to the usual text conditioning $c_{\theta}(y)$. To enable this, additional input channels are added to the first convolutional layer of $\boldsymbol{\epsilon_{\theta}}$ so that $z_{t}$ and $\mathcal{E}(c_{I})$ can be concatenated. Weights of these channels are initialized to zero, whereas rest of the model is initialized from the pre-trained Stable Diffusion 1.5 [49]. For architectural design choices, we resort to those made by Instruct-Pix2Pix [4].
@@
For training, we rely on a set $\mathcal{X}=\{(\mathbf{x}^{i},\mathbf{c}^{i},y^{i})_{i=1}^{N}\}$ of data triplets, each containing a clean image $\mathbf{x}_{i}$, an image with artifacts $\mathbf{c}_{i}$, and the corresponding edit instruction for fine-tuning $y_{i}$, to “teach” a denoiser to detect Gaussian artifacts and to generate a clean version of the conditioning image. For this, we build an artifact simulation engine comprising a full 3DGS model fitted to dense views, Sparse 3DGS fitted to few views, and camera interpolation and perturbation modules to use supervision of the full model at viewpoints beyond ground truth camera poses. The fine-tuning setup is illustrated in Fig. 2. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. For larger $M$, we observe that there are very few artifacts beyond standard Gaussian blur. To have diversity in edit instructions, we start with a base instruction - “Denoise the noisy image and remove all floaters and Gaussian artifacts.” and ask GPT-4 [1] to generate 50 synonymous instructions. During training, each noisy, clean image pair is randomly combined with one of them.
Figure 2: Artifact removal fine-tuning. Pairs of clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $36$ scenes. These are combined with one of 51 synonymous prompts generated by GPT-4 from a base instruction. SD $1.5$ is then fine-tuned with a dataset of $10.5K$ samples for the Gaussian artifact removal task.
Training the Denoiser
Using our synthetically curated dataset $\mathcal{X}$, we fine-tune SD 1.5 as follows:
@@
where $\mathbf{z}_{t}^{i}$ is the encoded image $\mathbf{x}^{i}$ diffused with sampled noise $\epsilon$ at time step $t$. Thus, the model is fine-tuned to generate clean images $\mathbf{x}^{i}$, conditioned on artifact images $\mathbf{x}^{i}$ and text prompts $y^{i}$.
@@ 4 Experiments
This section compares Sp2360 with state-of-the-art sparse-view reconstruction techniques. We also provide detailed ablation studies motivating specific design choices across different components of our approach. The setup for Iterative 3DGS (c.f. A.3) translates seamlessly to the iterative distillation procedure (c.f. A.4) with diffusion priors, and hence, we refrain from further analysis here.
@@ 4.1 Experimental Setup
We compare our approach against $7$ baselines. FreeNeRF [72], RegNeRF [42], DiffusioNeRF [70], and DNGaussian [32] are few-view regularization methods based on NeRFs or 3D Gaussians. ZeroNVS [52] is a recent generative approach for reconstructing a complete 3D scene from a single image. We use the ZeroNVS baseline introduced in ReconFusion [69], designed to adapt ZeroNVS to multi-view inputs. Conditioning the diffusion model on the input view closest to the sampled random view enables scene reconstruction for a general $M$-view setting. We also compare against 3DGS, the reconstruction pipeline for Sp2360, and against Sparse 3DGS, our self-created baseline. We cannot compare with ReconFusion [69] as neither their view splits nor code are publicly available, and their method builds on top of closed-source diffusion models. We expect to perform slightly worse due to weaker priors, but our approach is more lightweight and data-efficient.
@@ 4.2 Implementation Details
We implement our entire framework in Pytorch 1.12.1 and run all experiments on single $A100$ or $A40$ GPUs. We work with image resolutions in the 400-600 pixel range as this is closest to the output resolution of $512$ for both diffusion models. We set $\lambda_{1}=0.2$ (same as 3DGS) and vary $\lambda_{depth},\lambda_{pseudo}$ in $\{0.0,0.05,0.1\}$ for Sparse 3DGS. For in-painting, we set $\tau=0.8$ and fine-tune the in-painting module for 3000-5000 steps with LoRA modules of rank in $\{8,16,32\}$ (depending on $M$). We use a batch size of $16$ and learning rates of 2e-4 for the UNet and 4e-5 for the text encoder. The artifact removal module is trained for $2500$ iterations with batch size $16$ and learning rate 1e-4. To enable classifier-free guidance for both models, we randomly dropout conditioning inputs (text, mask, image, etc.) with probability of $0.1$ during training. The classifier-free guidance scales are set to $s_{I}=2.5$ and $s_{T}=7.0$. We use $t_{max}=0.99$ for both in-painting and artifact removal and linearly decrease $t_{min}$ for the in-painting module from $0.98$ to $0.90$, and from $0.98$ to $0.70$ for the artifact removal module. We sample both the in-painted and clean renders for $s=20$ DDIM sampling steps. We also linearly decay the weight of $\mathcal{L}_{sample}$ (c.f. A.4) from $1$ to $0.1$ over $30k$ iterations.
@@ 4.4 Ablation Studies
In Tab LABEL:tab:ablation and Fig 4, we ablate the relative importance of each component towards $360^{\circ}$ reconstruction. We pick the garden scene and its $9$ view split for this study. We first show benefits of the regularization heuristics in Sparse 3DGS over native 3DGS. We attempt reconstruction using either the in-painting or artifact removal module and show that their combination works best for both novel view synthesis and final reconstruction. Interestingly, leaving out the artifact removal module results in the best PSNR and SSIM scores across all variants, emphasizing that blurry reconstructions are rewarded by classical metrics, whereas FID and KID favor sharp, realistic details in novel views. We remove the iterative update formulation and add in-painted clean renders for all novel views simultaneously. This variant performs worse than Sparse 3DGS, highlighting the need for autoregressive scene generation in our setting. Additionally, we try to supervise the renderings at novel views using the original 3DGS objective. However, this performs slightly worse than Eq 11. Note that an SDS-based formulation [52] is not applicable here due to the two-step nature of generative view synthesis.
@@ Appendix A Appendix / Supplemental Material
In this section, we provide further details on certain heuristics and design choices of our pipeline not covered in the main paper. We begin with our strategy for creating $M$-view subsets of a scene (c.f A.1. We then provide details on the regularization introduced in our improved baseline to adapt 3DGS to sparse-view reconstruction (c.f A.2). We perform a toy experiment to determine the best choice of parameters for our iterative schedule (c.f. A.3). We end with some minor details on the distillation procedure to lift 2D diffusion priors to 3D Gaussians (c.f. A.4). Please also note our supplemental video, showing 360° reconstruction results.
@@ A.1 Creating M-view Sparse Sets
3DGS relies on COLMAP for a sparse point cloud initialization, and COLMAP requires sufficient feature correspondence across the input images to produce a 3D consistent sparse scene geometry. To produce a subset where the views are far apart to have sufficient scene coverage and at the same time having enough feature correspondence among the images to produce a 3D consistent (albeit sparse) point cloud, we devise a well-defined strategy such that view subsets across different scenes follow a certain heuristic and are not just randomly picked.
@@ A.2 Sparse 3DGS
Similar to [11], we observe that the periodic resetting of Gaussian opacities, designed to restrict over-reconstruction in a dense setting, harms generalization to novel views in a sparse setting. Either several Gaussians are removed through pruning or get trapped in local minima of the optimization process, not receiving enough updates to match the given scene. Visually, we observe much lower artifacts in novel views for extremely sparse settings ($M\leq 18$) if opacities $\alpha$ are optimized with no resets during training. The densification threshold for positional gradients - $\tau_{pos}$ is another hyperparameter that controls the overall number of Gaussians. Experimentally, we observe varying $\tau_{pos}$ between 2-5 times the default value for dense setting in 3DGS, leads to smoother geometry and prevents overfitting in a sparse setting for $M\leq 54$. We show visual evidence in Fig 8 and 9 below.
@@
Figure 9: Effect of densification threshold ($\tau_{pos}$) for $M=9$ in the $bicycle$ scene. A higher value of $\tau_{pos}$ in a view-constrained scenario limits artifacts and favours smoother geometry.
@@
We switch on $\mathcal{L}_{pseudo}$ after 2000 iterations when the optimization has stabilized, and Gaussians have learned at least a coarse scene geometry. We call this improved baseline Sparse 3DGS. Visually, we observe in Fig 10 that depth regularization through both ground truth and synthesized views helps 3DGS learn plausible geometry of the scene - like the structure of the bicycle and bench while smoothening out a lot of the grainy noise in the background.
@@
A.3 Iterative 3DGS

### v2

@@ 3.1 Sp2360 for Sparse-View 3D Reconstruction
For training 3DGS in an iterative fashion, as outlined above, special precaution has to be taken to prevent overfitting on the initial views. An optimal schedule involves finding the number of iterations per cycle $k$ and the hyperparameters for the original 3DGS [31]. We provide a detailed evaluation in App. A.4. In the following, we will detail the in-painting and artifact removal stages of our pipeline.
@@ 3.3 Removing Sparse-View Artifacts
We fine-tune an image-conditioned diffusion model [4] to edit images based on short, user-friendly edit instructions. Here, the UNet $\boldsymbol{\epsilon_{\theta}}$ is trained to predict noise in $z_{t}$ conditioned on a given image $c$, in addition to the usual text conditioning $c_{\theta}(y)$. To enable this, additional input channels are added to the first convolutional layer of $\boldsymbol{\epsilon_{\theta}}$ so that $z_{t}$ and $\mathcal{E}(c)$ can be concatenated. Weights of these channels are initialized to zero, whereas rest of the model is initialized from the pre-trained Stable Diffusion v1.5 checkpoint. For architectural design choices, we resort to those made by Instruct-Pix2Pix [4].
@@
For training, we rely on a set $\mathcal{X}=\{(\mathbf{x}^{i},\mathbf{c}^{i},y^{i})_{i=1}^{N}\}$ of data triplets, each containing a clean image $\mathbf{x}^{i}$, an image with artifacts $\mathbf{c}^{i}$, and the corresponding edit instruction for fine-tuning $y^{i}$, to “teach” a diffusion model how to detect Gaussian artifacts and generate a clean version of the conditioning image. For this, we build an artifact simulation engine comprising a 3DGS model fitted to dense views, Sparse 3DGS fitted to few views, and camera interpolation and perturbation modules to use supervision of the dense model at viewpoints beyond ground truth camera poses. The fine-tuning setup is illustrated in Fig. 2. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. For larger $M$, we observe that there are very few artifacts beyond standard Gaussian blur. To have diversity in edit instructions, we start with a base instruction - “Denoise the noisy image and remove all floaters and Gaussian artifacts.” and ask GPT-4 to generate 50 synonymous instructions. During training, each clean, artifact image pair is randomly combined with one of these 51 instructions.
Figure 2: Artifact removal fine-tuning. Pairs of clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $36$ scenes. These are combined with one of 51 synonymous prompts generated by GPT-4 [1] from a base instruction. SD v1.5 [51] is then fine-tuned with a dataset of $10.5K$ samples for the Gaussian artifact removal task.
Training the Artifact Removal Module
Using our synthetically curated dataset $\mathcal{X}$, we fine-tune SD v1.5 as follows:
@@
where $\mathbf{z}_{t}^{i}$ is the encoded image $\mathbf{x}^{i}$ diffused with sampled noise $\epsilon$ at time step $t$. Thus, the model is fine-tuned to generate clean images $\mathbf{x}^{i}$, conditioned on artifact images $\mathbf{c}^{i}$ and text prompts $y^{i}$.
@@ 4 Experiments
This section compares Sp2360 with state-of-the-art sparse-view reconstruction techniques. We also provide detailed ablation studies motivating specific design choices across different components of our approach. The setup for Iterative 3DGS (c.f. A.4) translates seamlessly to the iterative distillation procedure (c.f. A.5) with diffusion priors, and hence, we refrain from further analysis here.
@@ 4.1 Experimental Setup
We compare our approach against $7$ baselines. FreeNeRF [74], RegNeRF [44], DiffusioNeRF [72], and DNGaussian [33] are few-view regularization methods based on NeRFs or 3D Gaussians. ZeroNVS [54] is a recent generative approach for reconstructing a complete 3D scene from a single image. We use the ZeroNVS baseline introduced in ReconFusion [71], designed to adapt ZeroNVS to multi-view inputs. Conditioning the diffusion model on the input view closest to the sampled random view enables scene reconstruction for a general $M$-view setting. We also compare against 3DGS, the reconstruction pipeline for Sp2360, and against Sparse 3DGS, our self-created baseline. We cannot compare with ReconFusion [71] as their code is not publicly available, and their method builds on top of closed-source diffusion models. We expect to perform slightly worse due to weaker priors, but our approach is more lightweight and data-efficient.
@@ 4.2 Implementation Details
We implement our entire framework in Pytorch 1.12.1 and run all experiments on single $A100$ or $A40$ GPUs. We work with image resolutions in the 400-600 pixel range as this is closest to the output resolution of $512$ for both diffusion models. We set $\lambda_{1}=0.2$ (same as 3DGS) and vary $\lambda_{depth},\lambda_{pseudo}$ in $\{0.0,0.05,0.1\}$ for Sparse 3DGS. For in-painting, we set $\tau=0.8$ and fine-tune the in-painting module at $512\times 512$ resolution for 3000-5000 steps with LoRA modules of rank in $\{8,16,32\}$ (depending on $M$). We use a batch size of $16$ and learning rates of 2e-4 for the UNet and 4e-5 for the text encoder. The artifact removal module is trained at $256\times 256$ resolution for $2500$ iterations with batch size $16$ and learning rate 1e-4. To enable classifier-free guidance for both models, we randomly dropout conditioning inputs (text, mask, image, etc.) with probability of $0.1$ during training. The classifier-free guidance scales are set to $s_{I}=2.5$ and $s_{T}=7.0$. We use $t_{max}=0.99$ for both in-painting and artifact removal and linearly decrease $t_{min}$ for the in-painting module from $0.98$ to $0.90$, and from $0.98$ to $0.70$ for the artifact removal module. We sample both the in-painted and clean renders for $s=20$ DDIM sampling steps. We also linearly decay the weight of $\mathcal{L}_{sample}$ (c.f. A.5) from $1$ to $0.1$ over $30k$ iterations.
@@ 4.4 Ablation Studies
In Tab LABEL:tab:ablation and Fig 4, we ablate the relative importance of each component towards $360^{\circ}$ reconstruction. We pick the garden scene and its $9$ view split for this study. We first show benefits of the regularization heuristics in Sparse 3DGS over native 3DGS. We attempt reconstruction using either the in-painting or artifact removal module and show that their combination works best for both novel view synthesis and final reconstruction. Interestingly, leaving out the artifact removal module results in the best PSNR and SSIM scores across all variants, emphasizing that classical metrics reward blurry reconstructions, whereas FID and KID favor sharp, realistic details in novel views. We remove the iterative update formulation and add in-painted clean renders for all novel views simultaneously. This variant performs worse than Sparse 3DGS, highlighting the need for autoregressive scene generation in our setting. Additionally, we try to supervise the renderings at novel views using the original 3DGS objective. However, this performs slightly worse than Eq 12. Note that an SDS-based formulation [54] is not applicable here due to the two-step nature of generative view synthesis.
@@ Appendix A Appendix / Supplemental Material
In this section, we provide further details on certain heuristics and design choices of our pipeline not covered in the main paper. We begin with our strategy for creating $M$-view subsets of a scene (c.f A.1). We then provide details on the regularization introduced in our improved baseline to adapt 3DGS to sparse-view reconstruction (c.f A.2). Next, we compare our Gaussian artifact removal module with existing image-to-image diffusion models and show the importance of finetuned priors for adapting to this task (c.f A.3). We perform a toy experiment to determine the best choice of parameters for our iterative schedule (c.f. A.4). We end with some minor details on the distillation procedure to lift 2D diffusion priors to 3D Gaussians (c.f. A.5). Please also note our supplemental video, showing 360° reconstruction results.
@@ A.1 Creating M-view Sparse Sets
3DGS relies on COLMAP for a sparse point cloud initialization, and COLMAP requires sufficient feature correspondence across the input images to produce a 3D consistent sparse scene geometry. To produce a subset where the views are far apart to have sufficient scene coverage and at the same time have enough feature correspondence among the images to produce a 3D consistent (albeit sparse) point cloud, we devise a well-defined strategy such that view subsets across different scenes follow a certain heuristic and are not just randomly picked.
@@ A.2 Sparse 3DGS
Similar to [11], we observe that the periodic resetting of Gaussian opacities, designed to restrict over-reconstruction in a dense setting, harms generalization to novel views in a sparse setting. Either several Gaussians are removed through pruning or get trapped in local minima of the optimization process, not receiving enough updates to match the given scene. Visually, we observe much lower artifacts in novel views for extremely sparse settings ($M\leq 18$) if opacities $\alpha$ are optimized with no resets during training. The densification threshold for positional gradients - $\tau_{pos}$ is another hyperparameter that controls the overall number of Gaussians. Experimentally, we observe varying $\tau_{pos}$ between 2-5 times the default value for dense setting in 3DGS leads to smoother geometry and prevents overfitting in a sparse setting for $M\leq 54$. We show visual evidence in Fig 8 and 9 below.
@@
Figure 9: Effect of densification threshold ($\tau_{pos}$) for $M=9$ in the $bicycle$ scene. A higher value of $\tau_{pos}$ in a view-constrained scenario limits artifacts and favors smoother geometry.
@@
We switch on $\mathcal{L}_{pseudo}$ after 2000 iterations when the optimization has stabilized and Gaussians have learned at least a coarse scene geometry. We call this improved baseline Sparse 3DGS. Visually, we observe in Fig 10 that depth regularization through both ground truth and synthesized views helps 3DGS learn plausible geometry of the scene - like the structure of the bicycle and bench while smoothening out a lot of the grainy noise in the background.
@@
A.3 Baselines for Artifact Removal Module
In Fig 11, we compare the quality of diffusion samples from our proposed Artifact Removal Module (Sec 3.3) and two possible baseline solutions based on image-to-image diffusion. For this, we pick the $9$-view split of the bicycle scene and run our two-step view synthesis pipeline with the baselines. All $3$ methods take inpainted renders obtained from the Inpainting Module as input (Sec 3.2).
Render In-painted render SD Img2Img Instruct Pix2Pix Ours Ground Truth
View 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
View 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
View 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
View 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative comparison of Artifact Removal Module with image-to-image diffusion baselines. The baselines either produce unrealistic images with background collapse or fail to restore areas in the render corrupted by artifacts. Our finetuned prior, on the contrary, identifies Gaussian artifacts and removes them more efficiently to recover both foreground and background details in the scene.
Stable Diffusion Image-to-Image
Image-to-image diffusion is similar to text-to-image synthesis, but in addition to a text prompt, the final generation is also conditioned on an initial image. We use Stable Diffusion Image-to-Image [39] to try and remove Gaussian artifacts from inpainted renders. If $\mathbf{x}_{0}$ is the inpainted render, it is first encoded as $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$, and then noised corresponding to a certain timestep $t$ to obtain $\mathbf{z}_{t}$. The UNet $\boldsymbol{\epsilon_{\theta}}$ predicts the noise in $\mathbf{z}_{t}$ as:
$\hat{\epsilon_{t}}=\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,c_{\theta}(y)),$ (10)
Here $y=$“A photo of [V]”, where $[V]$ is a special token by textual inversion [17] from the $M$ reference images, to capture visual and semantic details of the scene. We set $t_{max}=0.3$ to ensure the output does not deviate too much from $\mathbf{x}_{0}$ and keeps at least geometry intact. However, the model has trouble differentiating Gaussian artifacts from image structures and hence cannot recover any scene details lost to artifacts. Instead, it creates images with oversmoothed geometry that look unrealistic and lack any high-frequency details of the original scene.
Instruct Pix2Pix
We try to obtain clean renders with the pretrained Instruct Pix2Pix (IP2P) [4] checkpoint using the same base instruction as our finetuned artifact removal module. However, IP2P was originally designed to perform image edits like replacing objects, changing the style or setting of a given image, etc, and not an image restoration task we aim to solve. As such, both the conditioning artifact image and the edit instruction are out-of-distribution for IP2P, and hence it fails to identify or remove any of the Gaussian artifacts.
Our finetuned artifact removal module, on the other hand, efficiently deals with blur, floaters, and color artifacts in $\mathbf{x}_{0}$ and recovers image structures lost to Gaussian artifacts. Artifacts induced at novel views by a 3D Gaussian representation are not observed in images used to train generic image-to-image diffusion models, necessitating the creation of our synthetically curated fine-tuning dataset.
A.4 Iterative 3DGS